### v3

# Reply with Sticker: New Dataset and Model for Sticker Retrieval

Abstract
Using stickers in online chatting is very prevalent on social media platforms, where the stickers used in the conversation can express someone’s intention/emotion/attitude in a vivid, tactful, and intuitive way. Existing sticker retrieval research typically retrieves stickers based on context and the current utterance delivered by the user. That is, the stickers serve as a supplement to the current utterance. However, in the real-world scenario, using stickers to express what we want to say rather than as a supplement to our words only is also important. Therefore, in this paper, we create a new dataset for sticker retrieval in conversation, called StickerInt, where stickers are used to reply to previous conversations or supplement our words [^1] We believe that the release of this dataset will provide a more complete paradigm than existing work for the research of sticker retrieval in the open-domain online conversation.. Based on the created dataset, we present a simple yet effective framework for sticker retrieval in conversation based on the learning of intention and the cross-modal relationships between conversation context and stickers, coined as Int-RA. Specifically, we first devise a knowledge-enhanced intention predictor to introduce the intention information into the conversation representations. Subsequently, a relation-aware sticker selector is devised to retrieve the response sticker via cross-modal relationships. Extensive experiments on the created dataset show that the proposed model achieves state-of-the-art performance in sticker retrieval [^2] The dataset and source code of this work are released at https://github.com/HITSZ-HLT/Int-RA..
Index Terms:
Sticker retrieval, Intention, Multi-modal learning, Conversation.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
With the rise of instant messaging applications, online chatting has become an essential part of daily life [1]. Stickers, as visual elements on social platforms, bring a dynamic and multifaceted dimension to conversations. Previous research on stickers has largely concentrated on sentiment analysis [2, 3, 4]. Due to their visual appeal, stickers uniquely contribute to fostering a lively and innovative conversational environment [5, 6]. Therefore, incorporating automatic sticker replies based on previous conversations into dialogue systems can significantly enhance the engagement and liveliness of interactions.
Figure 1: Two examples of stickers used in an online conversation.
Recent research endeavors of sticker retrieval have been dedicated to using stickers to supplement the current response, in order to strengthen the expression of emotion, attitude, or opinion [7, 8]. However, in the real-world scenario, we may also use stickers to reply to the previous conversation directly, rather than merely supplementing our words with stickers. One expected scenario of the sticker retrieval task is that suitable stickers can be retrieved for replies whether or not one has made a textual response. To illustrate our idea, we give examples shown in Figure 1. Figure 1 (a) shows the sticker retrieval scenario of using a sticker to supplement one’s current utterance. Figure 1 (b) is another sticker retrieval scenario, in which the user uses a sticker to reply to previous conversations directly. Therefore, in our work, we create a new sticker retrieval dataset to cover these two scenarios, called StickerInt, which is a comprehensive consideration for using stickers in social media conversation. Our StickerInt dataset contains 1,578 Chinese conversations with 12,644 utterances.
[…]
Our main contributions can be summarized as follows:
  • To facilitate the research of sticker retrieval, we create StickerInt, a novel and more comprehensive sticker retrieval dataset for sticker usage in conversation. Further, our StickerInt dataset provides an intention tag for each sticker towards the conversation, aiming at empowering models with more in-depth learning on sticker retrieval.
  • We propose a novel pipeline framework for sticker retrieval, in which a knowledge-enhanced intention predictor and relation-aware selector are devised to leverage the intention information and fine-grained sticker attributes to retrieve stickers for response.
  • Experiments conducted on our StickerInt dataset demonstrate that our proposed framework outperforms the baseline models.
II Related Work
II-A Sticker Dataset
Sticker analysis has attracted more and more attention in recent years. Numerous works fix their eyes on sticker-based multi-modal sentiment analysis and have proposed a wide variety of multi-modal sentiment analysis datasets [2, 3]. However, stickers, being visual images, demonstrate significant effectiveness within conversations beyond mere research into the stickers themselves. Consequently, several researchers have shifted their focus towards retrieving stickers based on the context of the conversation. [7] introduced a novel task termed meme incorporated open-domain conversation and further built a large dataset of 45k Chinese conversations with 606k utterances. A Chinese sticker-based multi-modal dataset for the sentiment analysis task (CSMSA) [4] was presented by collecting from eight public open chat groups with 28k text-sticker pairs and 1.5k annotated data. In these datasets, the stickers serve as a supplement to the current utterance. Nevertheless, in real-world scenarios, stickers are used not just as supplements to words but also as a means to directly express the user’s intentions. Therefore, we create a comprehensive dataset for sticker retrieval in conversation.
II-B Image Retrieval
Image retrieval aims to retrieve images that match a given query from a large collection of images. Early method [11] mainly relied on global feature-based image-text matching. Further, [12] proposed a Visual Semantic Reasoning Network (VSRN), which generates global features with regional semantic associations through a Graph Convolutional Network (GCN). To address the false negative problem in existing image-text matching benchmarks, [13] proposed an ITM framework integrating Language Guidance. In terms of improving retrieval efficiency and accuracy, [14] proposed a self-supervised fine-grained alignment module called SelfAlign. [15] proposed a feature-first approach to advance image-text retrieval by improving visual features. This method generates more comprehensive and robust visual features through a two-step interaction model and a multi-attention pooling module. Recently, with the development of large-scale pretraining models like CLIP [16] and ALIGN [17], significant performance improvements are achieved in image-text matching tasks by pretraining on massive vision-language data and then fine-tuning for specific downstream tasks. Different from the image retrieval methods, we focus on retrieving the image based on the conversation not just the text.
II-C Multi-modal Conversation
Several multi-modal studies have sought to improve the efficacy of conversational agents by enriching textual expressions of generated dialog responses through associative vision scenes [18, 19, 20, 21]. In contrast, sticker retrieval in the real-world scenario requires understanding the semantic expression and intention of context and the sticker image. Several works specifically paid attention to sticker features [22, 4]. [22] proposed a real-time system for sticker recommendation, which decomposes the sticker recommendation task into two steps including message prediction that the user is likely to send and an appropriate sticker substitution. Nowadays, using stickers for replies has become commonplace in social media interactions. There has been a growing number of works on sticker retrieval, which assists users in selecting the appropriate sticker for response. [8] proposed a sticker response selector model that employs a convolutional-based sticker image encoder and a self-attention-based multi-turn dialog encoder to obtain the representation of stickers and utterances. Then a deep interaction network is designed for obtaining deep matching and a fusion network is employed to output the final matching score. [7] presented a Meme incorporated Open-domain Dialogue (MOD) task and utilized pooling all sub-tasks like text generation and internet meme prediction into a unified sequence generation procedure to solve it. To tackle three challenges of inherent multimodality, significant inter-series variations, and complex multi-modal sentiment fusion. [4] integrated both the sticker text and series tag to holistically model sticker sentiment. They employed a flexible masked attention mechanism to selectively extract the most pertinent information crucial for the current sentiment assessment. These methods which directly match conversation context and stickers, overlook the expressive role of stickers in conversation, where emotions and intentions are conveyed in a visually engaging manner.
III StickerInt Dataset
This section introduces our new dataset StickerInt for sticker retrieval in detail. Specifically, Section III-A presents the construction of our StickerInt dataset, Section III-B introduces the annotation process of StickerInt, and Section III-C analyzes the created StickerInt by data statistics.
III-A Data Construction
We collected our dataset from a widely used social platform (WeChat [^3] https://weixin.qq.com/), which boasts a vast of conversations with stickers accessible for both individual and group chats. We select four open chat groups with active participants and collect their conversations. Among them, each group engages in open-domain online conversation, making the use of stickers more diverse. Note that we eliminate extraneous image elements like screenshots and photos.
We formulate stringent guidelines and policies for data preprocessing. To protect user privacy, users’ personal information (including real name, age, address, etc.) is deleted, and user IDs are anonymized in the data. Additionally, we exclude any content that may contain inappropriate, offensive, or insulting expressions. Furthermore, we segment the whole chat content within the dataset into multiple conversations to ensure the integrality and independence of each conversation. Based on this, we traverse each sticker in the chat history, capturing its associated context to derive conversations with stickers, which ensures that each sticker has a corresponding conversation context.
TABLE I: Statistics of StickerInt dataset. con. = conversation, Avg. = average.
Dataset Statistics Train Valid Test
# con. 1,269 155 154
# utterances 8,745 1,105 1,216
# tokens 47,895 5,950 5,778
# stickers 774 127 124
# users 48 41 40
Avg. utterances in a con. 6.89 7.13 7.90
Avg. users in a con. 2.91 2.71 2.78
Avg. tokens in an utterance 5.48 5.39 4.75
III-B Data Annotation
We recruited 5 experienced researchers with over 3 years of research experience in the field of multi-modal learning as annotators to check and label the golden sticker for each conversation, aiming at eliminating the impact of noise stickers on the research of sticker retrieval.
In addition, recognizing the sentiment, emotion, or feeling can help us better select stickers for replies. Therefore, considering the diversity and complexity of intentional expression in conversation, we get inspiration from [23] and use GoEmotions [24] to request annotators to supplement an intention tag for each sticker towards the conversation. We enlist the expertise of five annotators to label both the coarse and fine-grained intention of stickers based on the dialogues. Given that the outcome of the annotation process is closely tied to the annotators’ subjective judgment, each annotator is initially tasked with annotating 500 sticker samples. This initial task aids them in comprehending the annotation process and understanding the nuances of intention, thereby enhancing both the efficiency and accuracy of the subsequent labeling endeavors.
Figure 2: Visualization depicting the distribution of intention labels.
Figure 3: Similarity distribution among all stickers in our dataset. The x-axis represents the count of SSIM calculation, and the y-axis indicates the different range of sticker similarity.
III-C Dataset Analysis
The detailed statistics of our dataset are shown in Table I. In total, there are 1,578 conversations which contain 12,644 utterances, 59,623 tokens, and 1,025 stickers. Each conversation includes 8.01 utterances on average. The average number of users who participate in a conversation is 2.80. The average number of tokens in an utterance is 5.21. Furthermore, we also visualize the distribution of intention tags in Figure 2. We can see that the proportion of stickers used in different labels varies, which also demonstrates the diversity of stickers’ intention expressions in our dataset.
III-D Sticker Similarity
Stickers always share a similar style or contain the same cartoon characters. Intuitively, the more similar the stickers are, the more difficult it is to select the correct sticker from the sticker set. In other words, the similarity between stickers determines the difficulty of the sticker retrieval task. To investigate the difficulty of this task, we calculate the average similarity of all the stickers in our sticker set by the Structural Similarity Index (SSIM) metric [25]. We calculate the SSIM between each two stickers and normalize it into [0, 1]. The similarity distribution among our sticker data is shown in Figure 3, where the average similarity is 0.4016.
Figure 4: Illustration of the proposed Int-RA framework comprising intention-fused conversation context representation, attribute-aware sticker representation, and relation-aware sticker selector.
IV Methodology
In this section, we introduce our proposed Int-RA framework for sticker retrieval in detail. Assume that there is a conversation context $T=\{t_{1},...,t_{N_{t}},s_{j}\}$ and a sticker set $V=\{v_{1},...v_{N_{v}}\}$, where $t_{i}=(s_{i},u_{i})$, $s_{i}$ and $u_{i}$ represent the $i$-th user and utterance. $N_{t}$ and $N_{v}$ represent the number of utterances and stickers, respectively. The sticker retrieval task aims to select a suitable sticker from the sticker set $V$ based on the conversation context $T$ for the user $s_{j}$. In the $i$-th utterance $u_{i}=\{w_{1}^{i},...,w^{i}_{N_{w}^{i}}\}$, $w_{j}^{i}$ represents the $j$-th word in $u_{i}$, and $N_{x}^{i}$ represents the total number of words in $u_{i}$ utterance. Additionally, our dataset presents an annotated intention tag $y_{int}$ as supplemental information in each conversation $T$. In the dialog context $T$, $s_{i}$ represents a sticker image with a binary label $y_{i}$, indicating whether $s_{i}$ is an appropriate response for $T$. Each sticker with the intention label $m_{i}$ indicates the intention of the speaker to use the sticker to respond. Our goal is to learn a ranking model that can produce the correct ranking for each dialog.
Therefore, by leveraging the intention tags, we propose a pipeline framework (Int-RA) to deal with the sticker retrieval task. The architecture of our Int-RA is illustrated in Figure 4, which mainly comprises three components: 1) Intention-fused Conversation Context Representation, which introduces intention information to the learning of conversation context representation based on commonsense; 2) Attribute-aware Sticker Representation, which uses the multi-modal large language model (MLLM) to derive the representation of a sticker based on the attribute-based prompts; 3) Relation-aware Sticker Selector, which facilitates the retrieval of the relationships between conversation context and stickers by cross-modal attention.
IV-A Intention-fused Conversation Context Representation
For online chatting, we generally use stickers to express our sentiments, status, feelings, etc. Therefore, to explore the potential sentiment, status, or feeling expressed by the user towards a conversation context for sticker retrieval, we devise a knowledge-enhanced intention predictor, aiming to introduce intention information into the conversation context representation.
Inspired by recent works [26, 27], we adopt the commonsense knowledge base $\text{ATOMIC}_{20}^{20}$ [10], which contains knowledge not readily available in pre-trained language models and can generate accurate and representative knowledge for unseen entities and events. Specifically, we utilize the BART version of COMET [10] trained on this knowledge base to generate commonsense inferences of five relations including xIntent, xNeed, xWant, xEffect, xReact.
$C_{r}=\text{COMET}(T,r),$ (1)
[…]
$\textbf{{H}}^{T}=\text{M-BERT}(T\oplus C_{know}),$ (3)
Afterward, $\textbf{{H}}^{T}$ is input to a softmax classifier to infer the user’s intention $\hat{y}_{int}$ and optimized with cross-entropy loss, which can be described as:
$\rho=\text{softmax}(W\textbf{{H}}^{T}+b),$ (4)
$\hat{y}_{int}=\text{argmax}(\rho),$ (5)
$\mathcal{L}_{int}=-\sum_{i=1}^{n}y_{int}\log\rho.$ (6)
where $W$ and $b$ are the weight matrix and bias term, respectively. $n$ represents the number of samples, and $y_{int}$ is the ground-truth intention tag. In this way, we can obtain a deeper understanding of the intention that may be expressed towards the conversation context, and introduce intention information into the textual representation. The intention $\hat{y}_{int}$ is fed into M-BERT to obtain intention-fused textual representation $\textbf{{H}}^{Y}=\text{M-BERT}(\hat{y}_{int})$.
IV-B Attribute-aware Sticker Representation
Figure 5: Overview of attribute-aware sticker description generation.
To extract the key expression information and reduce unnecessary interference from irrelevant information, for the learning of stickers, we first devise four visual attributes, i.e. gesture $L_{G}$, posture $L_{P}$, facial expression $L_{F}$, and verbal $L_{V}$, to construct prompts for the strikers. Based on this, we use Qwen-VL [29] as the MLLM to produce attribute-aware sticker descriptions based on the above four attributes:
$\begin{split}&\{A_{G},A_{P},A_{F},A_{V}\}\\ &=\text{MLLM}(\{L_{G},L_{P},L_{F},L_{V}\}),\end{split}$ (7)
As demonstrated in Figure 5, we use several turns of interactions, including the system prompt like “This is a sticker used in conversation, please provide several keywords to describe the gesture/posture/facial expression/verbal.” to simulate the utterance generation ability of MLLM. Then, each attribute-aware sticker description is transformed into an attribute-aware sticker representation using M-BERT:
$\begin{split}&\textbf{{H}}^{A}=\{\textbf{{H}}^{A}_{G},\textbf{{H}}^{A}_{P},\textbf{{H}}^{A}_{F},\textbf{{H}}^{A}_{V}\}\\ &=\text{M-BERT}(\{A_{G},A_{P},A_{F},A_{V}\}).\end{split}$ (8)
Further, to learn the visual information of stickers, sticker $v_{i}$ first undergoes CLIP pre-trained ViT model [30] as a visual encoder to obtain visual representation $\textbf{{H}}^{I}$.
$\textbf{{H}}^{I}=\text{ViT}(v),$ (9)
Afterward, we adopt cross-modal attention between the visual representation and each attribute-aware sticker representation of the sticker to highlight the important regions in the sticker. In detail, we use two fully connected layers $f_{vis}$ and $f_{des}$ to project the visual representation and description representation into the same dimension $d$:
$\textbf{{h}}_{j}^{A}=f_{des}(\textbf{{H}}^{A}_{j}),\textbf{{h}}^{I}=f_{vis}(\textbf{{H}}^{I}),$ (10)
where $\textbf{{H}}^{A}_{j},j\in\{G,P,F,V\}$. $M_{j}\in\mathbb{R}$ indicates the relation between $\textbf{{h}}^{A}_{j}$ and the visual representation $\textbf{{h}}^{I}$, and can be expressed as:
$M_{j}=\text{softmax}(\frac{(\textbf{{h}}_{j}^{A}W^{Q})(\textbf{{h}}^{I}W^{K})^{\top}}{\sqrt{d_{k}}})(\textbf{{h}}^{I}W^{V}).$ (11)
where $W^{Q}\in\mathbb{R}^{d\times d_{q}},W^{K}\in\mathbb{R}^{d\times d_{k}},W^{V}\in\mathbb{R}^{d\times d_{v}}$ are randomly initialized projection matrices. We set $d_{k},d_{v},d_{q}=d/h$ for each of these parallel attention layers. $h$ is the number of heads in each multi-head attention layer.
Next, a max pooling operation is conducted on $M$, i.e., let $\mathcal{M}=\text{max}(M_{j})\in\mathbb{R}$ represent the relation score on the sticker by attribute-aware sticker descriptions. This attention learns to assign high weights to the important regions of the sticker that are closely related to each attribute-aware sticker description. We finally conduct a multiplication operation of each visual representation and relation score to obtain relation-aware visual representation $\textbf{{H}}^{R}$ for the sticker.
$\textbf{{H}}^{R}=\mathcal{M}\times\textbf{{h}}^{I}.$ (12)
IV-C Relation-aware Sticker Selector
Ultimately, we leverage the relation-aware sticker representations to perform cross-modal retrieval. We primarily implement the matching function using cosine similarity as cross-modal attention, which is defined as:
$CA=cos(\textbf{{H}}^{Y},\textbf{{H}}^{R}).$ (13)
We optimize our method to minimize a learning objective: $\mathcal{L}=\lambda_{1}\mathcal{L}_{ret}+\lambda_{2}\mathcal{L}_{int}$, where $L_{ret}$ is the loss for retrieval and $L_{int}$ for intention prediction. $\lambda_{1}$ and $\lambda_{2}$ are hyper-parameters that work as scale factors.
$\mathcal{L}_{ret}=\sum_{N}\max(\rho_{neg}-(1-\rho_{pos})+\text{margin}).$ (14)
where $\rho_{neg}$ and $\rho_{pos}$ correspond to the cosine similarity of non-true (negative) stickers and true (positive) stickers. The margin is the margin rescaling.
V Experiment
This section details the experimental settings and experimental results of our proposed Int-RA framework conducted on the created StickerInt dataset. We first present the experimental settings in Section V-A. Then, we introduce various compared methods in Section V-B. We analyze the performance of our approach through the Main Results (V-C), Ablation Study (V-D), Effect of Number of Utterances (V-F), and Effect of Different Attributes (V-G). Finally, to more intuitively demonstrate the performance of our method, we visualize the relations in Section V-I and present some interactive cases in Section V-H.
V-A Experimental Settings
Implement details. We adopt Multi-lingual BERT [28] as the text encoder to derive the textual representation. The CLIP pre-trained ViT model [30] is employed as the image encoder to derive the visual representation. we adopt Qwen-VL [29] to generate sticker descriptions  [^4] In the preliminary experiment, we also try other MLLMs such as mini-GPT4 [31] and Llava [32]. We found that Qwen-VL performed slightly better.. We set the batch size to 4 and use Adam optimizer [33] as our optimizing algorithm. The learning rate is set to $1\times 10^{-4}$. Both $\lambda_{1}$ and $\lambda_{2}$ are set to 1.
Evaluation metrics. Three widely used evaluation metrics are applied in our experiments: mean of average precision (mAP), top N-precision (P@N).mAP is a widely accepted criterion for assessing retrieval accuracy [34]. P@N evaluates the precision of the top N predictions. Here, we mainly present the results for P@1, P@3, and P@5. Moreover, the PR curve visually illustrates the trade-off between precision and recall at various thresholds. Notably, if the retrieved sticker matches the intention label of the ground truth sticker, we consider the result correct, as multiple stickers can serve as responses to the same conversation.
V-B Compared Methods
To evaluate the performance of our model, we compare the proposed Int-RA with several baseline methods, including existing sticker retrieval methods, recent text-to-image retrieval approaches, and large language models (LLMs).
  • Sticker retrieval methods: MOD [35], which leverages a unified generation network to produce multi-modal responses. SRS [36] which learns sticker representations and utterance context in the multi-turn dialog.
  • Text-to-image retrieval methods: LGUR [37] which enhances feature granularity alignment and performance through a transformer-based approach. IRRA [38] which leverages implicit relation reasoning and similarity distribution matching for improved cross-modal alignment. PCME [39] which presents a probabilistic embedding model for cross-modal retrieval. CLIP [16] which introduces a vision model trained on internet image-text pairs.
  • Large language models: Baichuan2, LLama3, ChatGLM3, Qwen2, Qwen-VL, and LLaVA.
TABLE II: Experimental results (%) of various methods in sticker retrieval and text-to-image retrieval methods. Bold indicates that our method surpasses other models. We assert significance if p-value < 0.05 under a t-test with the baseline models. w/o means without.
Model P@1 P@3 P@5 mAP
Sticker retrieval methods
SRS 1.30 3.25 6.49 3.66
MOD 5.84 9.74 14.29 9.72
[…]
IRRA 1.95 6.49 9.74 6.10
PCME 3.90 11.69 19.48 11.40
LGUR 9.09 11.69 15.58 12.10
CLIP 5.19 12.34 20.13 12.89
[…]
LLaVA 6.49 9.74 14.94 34.42
Qwen-VL 9.09 24.68 30.52 33.41
Qwen2 13.64 25.32 27.92 35.27
Baichuan2 14.29 24.68 26.62 34.94
Llamma3 15.58 20.78 27.27 39.76
ChatGLM3 15.58 25.32 31.17 36.29
[…]
Int-RA (ours) 18.18 37.66 40.91 53.37
w/o attribute 17.53 35.06 46.75 44.85
w/o intention 17.53 35.71 41.56 43.75
w/o knowledge 18.18 29.22 39.61 39.36
TABLE III: Experimental results (%) of different LLMs with or without intention.
Model Sticker Response Utterance Response
P@1 P@3 P@5 mAP P@1 P@3 P@5 mAP
LLaVA 6.49 9.74 14.94 34.44 16.23 20.78 29.87 39.35
-w/ intention 8.44 25.97 27.27 33.71 11.69 19.48 29.22 35.34
[…]
-w/ intention 9.09 25.97 32.47 37.39 9.09 23.38 28.57 39.85
[…]
-w/ intention 12.34 25.32 28.57 31.52 12.34 21.43 27.92 36.17
[…]
-w/ intention 12.99 22.08 25.97 35.48 12.34 25.97 28.57 33.84
[…]
-w/ intention 10.39 22.73 25.97 33.50 9.74 18.83 24.03 33.29
[…]
-w/ intention 11.04 20.13 25.32 35.50 11.69 25.97 32.47 37.69
V-C Main Results
We examine the performance of our Int-RAframework in comparison with baselines across each evaluation metric and report the results in Table V-B. We also assess the significance of performance differences between the two runs using a two-tailed paired t-test, with strong significance at $\alpha=0.01$ denoted by . It can be observed that our Int-RA consistently outperforms all baselines, demonstrating the effectiveness of the proposed Int-RA in sticker retrieval. We can also notice an improvement in results as the value of N increases in Top N-precision, as more results can be utilized to expand the scope of potential matches with relevant labels.
For sticker retrieval methods, MOD and SRS perform significantly poorer compared to our Int-RA. This further highlights the efficacy of our approach in first predicting intention before conducting matching, emphasizing the crucial role of intention as a bridging component in the process.
For text-to-image retrieval methods, they prioritize capturing semantic relationships between text and image content. However, since they are not explicitly designed for sticker retrieval scenarios, exhibiting inferior performance compared to our framework. In addition, an interesting observation is the overall superior performance of text-to-image retrieval methods compared to sticker retrieval baseline methods. This disparity can be attributed in part to the model design of SRS and MOD, where both models are devised to retrieve suitable stickers from a limited set of similar sticker candidates. Consequently, more attention is devoted to distinguishing local information among similar sticker expressions. In contrast, our dataset encompasses a more diverse range of stickers in real-world conversations, requiring the model to analyze and recognize more complex visual features. This demonstrates the advantage of the new dataset and the proposed framework in this work.
For Large language models, we design a prompt that integrates the current session to generate the sticker description for each session leveraging the in-context learning ability of LLMs [40, 41]. We then retrieve the appropriate sticker based on the generated sticker description and the combination of intent and sticker attributes using OpenAI’s LLM-based embedding model (text-embedding-ada-003) [42]. We can observe that LLM models perform better than sticker retrieval and text-to-image retrieval methods. This superior performance is attributed to the extensive parameterization and complex network architecture of LLMs, which greatly enhance their ability to understand and generate complex language and image descriptions. For P@1, P@3, and P@5, ChatGLM3, and Llama3 perform the best, while for mAP, Llama3 achieves the highest performance, reaching 39.76%. Compared to our IGSR method, all baseline models perform significantly worse. This further emphasizes the effectiveness of our approach in intent derivation, highlighting the critical role of intent as a bridging component in the process.
V-D Ablation Study
We also conduct an ablation study on the use of knowledge and attributes. The evaluation results are shown in Table V-B. The performances of all ablation models are worse than those of the complete model under all metrics, which demonstrates the necessity of each component in our approach. Note that the removal of attributes ("w/o attribute") results in considerable performance degradation, indicating that utilizing attributes can make better learning of sticker representation in different sticker properties. Notably, by observing the performance of w/o attribute on P@5, we find that the impact of attributes is not as significant when a larger number of stickers are recalled. In addition, the removal of commonsense knowledge ( "w/o knowledge" ) sharply degrades performance, which verifies the importance of knowledge in understanding conversation context. It is worth noting that the absence of the intention tag ("w/o intention") leads to a more significant decline in performance, with the mAP and P@5 scores decreasing by 0.62% and 14.01%, respectively. This demonstrates the crucial role of the intention tag in improving the accuracy and relevance of sticker retrieval tasks, highlighting the importance of considering users’ intentions in retrieving appropriate stickers.
V-E Effect of Using Intention
To analyze the impact of intention, We introduce the intention into various LLMs as mentioned in the baseline. The results are shown in Table V-B. "w/ intention" indicates that the large language model first predicts the intent of the response and then generates the reply based on the previous conversation context and the intent. "Sticker Response" and "Utterance Response" indicate whether the model generates a description of the sticker or generates a response for the user. Overall, the sticker description approach performs better than the text response approach, indicating that generating a sticker description can more directly highlight the key points of the response, thus improving the retrieval of the corresponding sticker. Furthermore, we can also find that predicting the intention first and then generating the response is less effective than directly generating the response for most LLMs. That is to say, an incorrect initial prediction of intention can lead to inappropriate responses, ultimately reducing performance. This suggests that merely introducing intention does not guarantee improved model performance. In contrast, our approach enables the model to simultaneously perform sticker retrieval and learn response intent, resulting in superior performance.
Figure 6: Performance of our approach on all metrics with different numbers of utterances.
V-F Effect of Number of Utterances
To examine and analyze the impact of the number of utterances over the performance of our proposed Int-RA framework, we conduct experiments by varying the number from 2 to 10 and demonstrate the results in Figure 6. We observe a similar trend across all evaluation metrics: mAP, P@1, P@3, and P@5. The results initially increase until the number of utterances reaches 6, after which they decrease as the number of utterances continues to increase. Two potential reasons may explain this phenomenon. Firstly, in limited contexts, the model can effectively capture features, resulting in improved performance as the amount of information increases. Secondly, the utility of utterance context may play a role. Utterances appearing too early before the sticker response may be irrelevant to the sticker and introduce unnecessary noise. In this dataset, it appears that 6 utterances are optimal.
TABLE IV: Experimental results (%) of different attributes. $\surd$ represents the used attribute. Ges., Pos., Face., and Ver. indicate gesture, posture, facial expression, and verbal, respectively.
Ges. Pos. Face. Ver. P@1 P@3 P@5 mAP
$\surd$ 9.09 22.08 31.17 38.32
$\surd$ 10.39 25.97 37.66 49.72
$\surd$ 9.09 27.27 37.66 39.76
$\surd$ 12.99 26.62 37.01 48.97
[…]
$\surd$ $\surd$ 11.04 35.06 36.36 39.12
$\surd$ $\surd$ 10.39 30.52 42.86 46.28
$\surd$ $\surd$ 12.99 29.22 42.21 44.82
$\surd$ $\surd$ 14.94 27.92 33.77 45.98
$\surd$ $\surd$ 13.64 25.32 37.66 49.95
[…]
$\surd$ $\surd$ $\surd$ 15.58 29.22 40.26 48.66
$\surd$ $\surd$ $\surd$ 14.29 25.97 38.31 46.00
[…]
V-G Effect of Using Different Attributes
In the process of attribute-aware sticker representation, four visual attributes are utilized in our proposed Int-RA to represent the key expression information of the sticker. This section examines the effectiveness of different attributes. The results of various scenarios of attribute combinations are shown in Table IV. It can be seen that the performance overall improves with the increase in the number of attributes used, using all the attributes achieves the best performance. Using only one attribute significantly degrades the performance, indicating that the visual information can not be learned sufficiently from a single perspective. That is, relying solely on this single attribute is insufficient for capturing the full expressive range necessary for accurate sticker representation. Further, incorporating multiple attributes provides richer information, thereby leading to improved performance. This concludes that a more holistic approach that combines multiple attributes to understand the visual information of stickers is essential for optimal performance.
Figure 7: Examples of conversation context and top-3 stickers retrieved by our method.
V-H Case Study
Several interactive cases retried by our approach are provided in Figure 7. These conversation samples suggest that our pipeline framework holds the capacity to provide sticker-incorporated expressive communication. From examples (a) and (b), we can observe that our approach tends to favor stickers with similar actions and facial expressions, with the characteristics of emoji stickers often being manifested through detailed features such as gestures and facial expressions. This demonstrates the effectiveness of using visual attributes to enhance the learning of stickers. However, in example (c), the final prediction is incorrect, likely due to the diverse styles of stickers, which remains a challenge in the current state of sticker recognition.
Furthermore, in example (a), User 9’s response is not a direct reply to the preceding utterance but rather addresses the emotional expression in the historical conversation "I lost my campus card". This shows the difficulty of understanding conversation context in multi-user conversation, resulting in the increased challenge of sticker retrieval. Consequently, in future research, user information can be considered to further improve the performance of sticker retrieval.
Figure 8: Visualization of the relation score on stickers. The darker the color, the higher the relation score.
V-I Visualization
To analyze how our Int-RA learns the important information about stickers, we visualize the relation score $\mathcal{M}$ (Equation 12) of three stickers in Figure 8. For example (a), where the character appears very angry. This indicates that the representation of this sticker heavily relies on this facial expression. Our Int-RA can effectively catch the important information in the sticker by the relation score placement on the character’s face. Moreover, the relation score also attends to the character’s gestures. For instance, in Case (b), where the character is depicted as holding chopsticks with one hand and supporting the face with the other, we observe attention focused on his hand, suggesting that our Int-RA learns key points of body language. Furthermore, considering that the relation score comprehensively considers four properties of stickers, as illustrated in Case (c), we observe that our Int-RA pays attention to both facial expressions and gestures simultaneously, thereby learning accurate visual information about the sticker.
VI Conclusion and Future Work
We create a new dataset for sticker retrieval, called StickerInt. Unlike previous studies that view stickers merely as a supplement to the current utterance, our new dataset can cover two real-world scenarios of using stickers in online conversation: using stickers to reply to previous conversations or supplement our words. Based on the new dataset, we propose Int-RA, a framework for sticker retrieval in conversation. In which, the intention information is leveraged in the learning of conversation context. Further, we devise four novel visual attributes, i.e. gesture, posture, facial expression, and verbal, to improve the learning of stickers. Based on this, a relation-aware sticker selector is explored to retrieve the sticker for the conversation. Extensive experiments conducted on our StickerInt dataset demonstrate that our proposed approach achieves outstanding performance in sticker retrieval.
In the future, we will focus on developing advanced models that can better handle the diversity of sticker styles in real-world conversations and improve the accuracy of sticker retrieval in multi-user, multi-turn dialogues. Furthermore, we plan to explore personalizing sticker recommendations and exploring cross-cultural differences in sticker usage to enhance the versatility and applicability of our models across various user groups and scenarios.
VII Limitations
The limitations of this work are mainly twofold. Firstly, stickers have diverse styles ( e.g. cartoon, animal, etc.) in real-world conversations, which might affect the performance of the sticker retrieval task. Additionally, real-world conversations often involve multiple users engaging in multi-turn conversations. In such scenarios, our method may not fully capture the complexities of interactions among multiple users. Future research could focus on addressing these limitations by exploring more sophisticated models or incorporating additional contextual information to improve the performance of the sticker retrieval task.
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (62176076), the Natural Science Foundation of Guangdong (2023A1515012922), the Shenzhen Foundational Research Funding (JCYJ20220818102415032), and the Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies (2022B1212010005).